OpenStudy (princeevee):

Whats (-16) - (-40) - 7

TheSmartOne (thesmartone):

Hint: a - (-b) = a + b

OpenStudy (princeevee):

so it equals 9?

TheSmartOne (thesmartone):

show me your work :D

OpenStudy (princeevee):

(-16) - (-40) = -16 - 7?

TheSmartOne (thesmartone):

how is that possible you can't replace -40 with 7

TheSmartOne (thesmartone):

ok, first, what is -16 - (-40) = ?? use the hint which I gave you

OpenStudy (princeevee):

it equals 16? Brb

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne Hint: a - (-b) = a + b \(\color{blue}{\text{End of Quote}}\) use this

OpenStudy (princeevee):

so (-16) - (-40) = 16 - 7 = 9?

OpenStudy (princeevee):

@TheSmartOne u there?

TheSmartOne (thesmartone):

nooooooooo :/

pooja195 (pooja195):

(-16) - (-40) - 7 remove the parentheses -16 - -40 - 7 -+-=+ -16+40-7

pooja195 (pooja195):

Can you solve that?

OpenStudy (princeevee):

uh. i cant really, its hard for me.

TheSmartOne (thesmartone):

You don't own a calculator? :P

pooja195 (pooja195):

ok 40-16=?

TheSmartOne (thesmartone):

Here is another hint: -16+40-7 = 40 - 16 - 7 You just need to solve 40 - 16 - 7

OpenStudy (alexandervonhumboldt2):

it is a problem which can be solved without a calculator

OpenStudy (princeevee):

17

TheSmartOne (thesmartone):

Correct! :)

OpenStudy (alexandervonhumboldt2):

yeah 17

pooja195 (pooja195):

Well done =)